OpenStudy (patrickweewoos):

Derrick and Mark are brothers who live in the same house. Mark is home, and Derrick is 5 mi from home. Both boys begin riding their bikes at the same time. Derrick rides directly home at a constant rate of 15 mph, and Mark rides away from home at a constant rate of 18 mph. Let d represent distance from home, and let t represent time in hours. Which system models this situation?

OpenStudy (gorv):

situation when they will meat ??

OpenStudy (gorv):

distance travelled by both must be equal to 5 mile

OpenStudy (gorv):

distance= speed *time

OpenStudy (patrickweewoos):

so speed x"s time

OpenStudy (gorv):

derrick with 15mph travel in time =??

OpenStudy (cwrw238):

their combined speed is 33 mph

OpenStudy (cwrw238):

- assuming their are approaching each other directly

OpenStudy (patrickweewoos):

Those are the answer choices

OpenStudy (gorv):

derrick speed=15 time=t distance=speed * time

OpenStudy (gorv):

mark speed=18mph time=t distance=??

OpenStudy (gorv):

sum of distance travelled by both =5 mi

OpenStudy (patrickweewoos):

A. d = 5 − 15t d = 18t B. d = 15 − 5t d = 18t C. d = 5 + 15t d = 18t D. d = 5 − 15t d = 5 + 18t

OpenStudy (patrickweewoos):

The answer choices

OpenStudy (patrickweewoos):

@perl can you help

OpenStudy (patrickweewoos):

@gorv

OpenStudy (gorv):

for derrick

OpenStudy (gorv):

speed= 15 time=t distance=speed*time so distance =?? plug in the values and tell what u got??

OpenStudy (patrickweewoos):

38

OpenStudy (patrickweewoos):

i dont understand

OpenStudy (gorv):

how u got 38

OpenStudy (gorv):

distance= speed*time what is speed of derrick

OpenStudy (patrickweewoos):

15+18=33 so i added 5 heres the answer choices, A. d = 5 − 15t d = 18t B. d = 15 − 5t d = 18t C. d = 5 + 15t d = 18t D. d = 5 − 15t d = 5 + 18t

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (patrickweewoos):

can you please just help me @gorv

OpenStudy (perl):

I can help

OpenStudy (gorv):

derrick speed=??

OpenStudy (patrickweewoos):

ok @perl do you know this I have to go somewhere in a few and have 5 math questions plz help

OpenStudy (gorv):

@perl lol who stopped u

OpenStudy (perl):

Derrick and Mark are brothers who live in the same house. Mark is home, and Derrick is 5 mi from home. Both boys begin riding their bikes at the same time. Derrick rides directly home at a constant rate of 15 mph, and Mark rides away from home at a constant rate of 18 mph. Let d represent distance from home, and let t represent time in hours.

OpenStudy (patrickweewoos):

Look guys i just need help

OpenStudy (perl):

I can help

OpenStudy (perl):

one sec

OpenStudy (patrickweewoos):

A. d = 5 − 15t d = 18t B. d = 15 − 5t d = 18t C. d = 5 + 15t d = 18t D. d = 5 − 15t d = 5 + 18t theres my choices for answers

OpenStudy (perl):

i will draw the situation

OpenStudy (patrickweewoos):

Ok :D

OpenStudy (perl):